OpenStudy (anonymous):

The endpoints of a line segment graphed on a Cartesian coordinate system are (2, -5) and (-4, 2). What are the coordinates of the midpoint of the segment?

OpenStudy (anonymous):

A. (-4, 2) B. (2, -3.5) C. (-1, 1.5) D. (1, -3.5) E. (-1, -1.5) F. (3.5, 2)

OpenStudy (anonymous):

The midpoint of a segment is the average of the x-values, and the average of the y-values - it's in the middle of the two points. The average of the x-values is -1, and the average of the y-values is-1.5. So choice C would be correct. If you don't understand how I averaged the numbers, I can explain that.

OpenStudy (anonymous):

i dont can you explain?

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

E

OpenStudy (whpalmer4):

@HarryPotter5777 had the right idea, he just made a sign error. x coordinate is average of 2, -4 add 2 and -4: 2 + -4 = 2-4 = -2 divide by 2: -2/2 = -1 y coordinate is average of -5, 2 add -5, 2: -5 + 2 = -3 divide by 2: -3/2 = -1.5 midpoint is at (-1,-1.5)